# SpatialGen: Layout-guided 3D Indoor Scene Generation

### v3

@@ Abstract
Creating high-fidelity 3D models of indoor environments is essential for applications in design, virtual reality, and robotics. However, manual 3D modeling remains time-consuming and labor-intensive. While recent advances in generative AI have enabled automated scene synthesis, existing methods often face challenges in balancing visual quality, diversity, semantic consistency, and user control. A major bottleneck is the lack of a large-scale, high-quality dataset tailored to this task. To address this gap, we introduce a comprehensive synthetic dataset, featuring 12,328 structured annotated scenes with 57,440 rooms, and 4.7M photorealistic 2D renderings. Leveraging this dataset, we present SpatialGen, a novel multi-view multi-modal diffusion model that generates realistic and semantically consistent 3D indoor scenes. Given a 3D layout and a reference image (derived from a text prompt), our model synthesizes appearance (color image), geometry (scene coordinate map), and semantic (semantic segmentation map) from arbitrary viewpoints, while preserving spatial consistency across modalities. SpatialGen consistently generates superior results to previous methods in our experiments. We are open-sourcing our data and models to empower the community and advance the field of indoor scene understanding and generation.
@@ 1 Introduction
A major consideration in developing 3D scene generation methods is the trade-off between realism and scene diversity. Procedural modeling methods [33, 54, 9] leverage hand-crafted heuristic rules and geometric constraints in the graphics engines, which produce highly realistic and physically plausible indoor environments. However, these scenes lack diversity. Recent 3D generative methods automatically generate scene layouts [28, 45] or other 3D representations like NeRFs [1] and 3D Gaussians [21]. But these methods exhibit limited layout and appearance realism, primarily due to the scarcity of annotated 3D data. In comparison, image-based methods utilize diffusion models to generate panoramas [46, 38] or multi-view images [11, 44] followed by 3D reconstruction. By leveraging powerful 2D priors, these methods show promise in striking a better balance between realism and scene diversity. Image-based methods, however, face additional challenges in multi-view semantic consistency. While recent video generation methods [55, 26] have improved temporal coherence, synthesizing semantically consistent content when exploring beyond input views remain highly challenging.
@@
To overcome these limitations, we collect a new indoor scene dataset on a much larger scale. Our dataset features 4.7M panoramic images with precise 2D and 3D layout annotations, spanning 57,440 rooms and 12,328 scenes. With this dataset, we take a new approach to 3D scene synthesis by building a scalable multi-view diffusion (MVD) model conditioned on 3D layout priors, which achieves high semantic consistency while maintaining the realism and scene diversity in the results.
@@
  • We introduce a new large-scale dataset featuring over 4.7M panoramic images of 57,440 rooms and precise 2D and 3D layout annotations. This dataset fills a critical gap in 3D scene modeling by providing comprehensive multi-view data with structural annotations.
@@ 3 SpatialGen Dataset
As shown in Figure 2, we create physically plausible camera trajectories that navigate smoothly through each scene while avoiding obstacles. These trajectories are sampled at $0.5\textrm{m}$ intervals to ensure comprehensive spatial coverage. For each viewpoint, we generate photorealistic panoramic renderings using an industry-leading rendering engine, capturing color, depth, normal, semantic, and instance segmentation data. We further convert the panoramic image into multiple perspective images using equilib [14]. To ensure both quality and diversity, we apply rigorous filtering criteria during dataset curation, resulting in 12,328 distinct scenes encompassing 57,440 individual rooms with diverse room types. Each scene is annotated with precise 3D layouts and divided into 57,390/50 scenes for training/testing, respectively.
Figure 2 also shows some panoramas and 3D layout annotation from our dataset. Our dataset offers comprehensive structural layout annotations, including architecture elements (i.e., walls, doors, and windows). We further simulate diverse camera motion patterns from panoramic video data to train and evaluate the generation capabilities of existing approaches – a crucial advantage over limited rule-based trajectories from existing dataset. Additional dataset samples are provided in the appendix.
@@ 5.2.1 Comparison to Score Distillation Methods
As one can see, our method performs slightly worse than SceneCraft when trained solely on Hypersim. We hypothesize that Hypersim is too small for powerful latent multi-view diffusion models like the one employed by our method. When trained on both Hypersim and our datasets, our method outperforms both SDS methods on all metrics. Notably, SpatialGen achieves a significantly higher image-reward score than models trained solely on Hypersim, validating the benefit of our large-scale dataset for high-quality 3D scene generation. Furthermore, when tested on our dataset, SpatialGen consistently outperforms the baselines, with a significantly higher image-reward score.
@@
The advantage of our method can be better observed in Figure 6. On both Hypersim and our datasets, SpatialGen generates photorealistic scenes with superior details that are well-aligned with the specified layout. In contrast, SceneCraft struggles to balance layout adherence with text-prompt fidelity, and Set-the-Scene takes nearly two hours to synthesize a radiance field that still lacks fine-grained details.
@@ 5.2.2 Comparison to Panorama-as-Proxy Methods
Next, we compare our method to panorama-as-proxy methods. We choose Ctrl-Room [8] as the baseline because the source code and checkpoints are available.
We conduct experiments on both Structured3D and our datasets. For Structured3D dataset, we use ground-truth layouts to ensure consistent inputs for both Ctrl-Room and our method. For our dataset, we again use the Inward Orbit camera trajectory for consistency.
@@
5.3 Image-to-3D Scene Generation
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
Forward
SpatialGen (w/o layout) 11.47 0.49 0.59 67.96
SpatialGen (w/ layout) 17.59 0.69 0.32 34.98
Inward orbit
SpatialGen (w/o layout) 12.57 0.48 0.54 64.14
SpatialGen (w/ layout) 17.30 0.66 0.33 35.57
Outward orbit
SpatialGen (w/o layout) 11.14 0.60 0.47 76.73
SpatialGen (w/ layout) 13.32 0.59 0.46 57.76
Random walk
SpatialGen (w/o layout) 11.26 0.45 0.59 98.42
SpatialGen (w/ layout) 14.07 0.62 0.45 52.10
Quantitative Results
Qualitative Results
@@ 6 Conclusion & Limitations
We present SpatialGen, a novel framework for layout-guided 3D indoor scene synthesis. At the core of our pipeline is a multi-view multi-modal diffusion model, which generates images with high visual quality and geometric consistency. To train this model, we collect a new synthetic indoor scene dataset with 4.7M panoramic renderings of 57,440 rooms and the 3D layout annotations. These advancements open new possibilities for downstream applications such as interior design, embodied AI, and virtual/augmented reality.
@@
First, the cross-view and cross-modal attention introduces additional computational cost to the multi-view diffusion model, which limits SpatialGen to generate a relatively small number of images at a time. Moreover, the camera sampling strategy might affect the generation quality. We plan to address these challenges in the future.
@@ Acknowledgments
This work was partially supported by the Key R&D Program of Zhejiang Province (2025C01001) and the HKUST project 24251090T019. We would like to thank the engineering team in Manycore Tech Inc., – Yingqi Shen, Liangbin Hu, and Fuchun Dong – for their exceptional effort in supporting building the large-scale SpatialGen dataset. We thank Chenfeng Hou for experiments on the SDS-based method, Zhiwei Wang for assistance with ControlNet testing, and Kunming Luo for valuable suggestions regarding the figures.
@@ A.1 Dataset Construction
The final dataset consists of 12,328 distinct scenes, 57,440 individual rooms covering a variety of room types, and 4.7M photo-realistic panoramic renderings. The total floor area across all scenes is approximately $914,687\textrm{m}^{2}$.
@@ A.3 Dataset Visualization
Appendix B Additional Results
In this section, we show more results of Text-to-3D Scene and Image-to-3D Scene generation.
@@ Appendix B Additional Results
Text to 3D scene generation
@@
Image to 3D scene generation
Building upon the large-scale SpatialGen dataset, we evaluate our method across four diverse camera trajectories. As shown in Figure 18, given a 3D layout and a reference image (highlighted in orange box), our method successfully generates 3D-consistent novel views and synthesizes semantically plausible content for areas beyond the original input view.
Appendix C Additional Experiments
In this section, we conduct additional ablation studies Section C.1. Furthermore, we show some generation results from a common video (unposed) captured in indoor scenes, using the state-of-the-art scene understanding model SpatialLM [27] in Section C.2.
C.1 Ablation Studies
@@
C.2 Creating New Scenes from Video
To demonstrate the versatility of SpatialGen, we apply it to the challenging task of generating novel 3D scenes from videos. By leveraging a state-of-the-art layout estimation model, SpatialLM [27], we get the reconstructed 3D layout from the video. Then, we perform text-to-3D scene generation conditioned on this layout and additional user-provided text prompts. This approach allows us to generate entirely new scenes that preserve the structural layout of the original video while altering its stylistic and semantic content based on the text description. We validate this video-to-new-scenes application on the SpatialLM test set. Figure 21 shows qualitative results.

### v4

@@ Abstract
Creating high-fidelity 3D models of indoor environments is essential for applications in design, virtual reality, and robotics. However, manual 3D modeling remains time-consuming and labor-intensive. While recent advances in generative AI have enabled automated scene synthesis, existing methods often face challenges in balancing visual quality, diversity, semantic consistency, and user control. A major bottleneck is the lack of a large-scale, high-quality dataset tailored to this task. To address this gap, we introduce a comprehensive synthetic dataset, featuring 12,328 structured annotated scenes with 57,431 rooms, and 4.7M photorealistic 2D renderings. Leveraging this dataset, we present SpatialGen, a novel multi-view multi-modal diffusion model that generates realistic and semantically consistent 3D indoor scenes. Given a 3D layout and a reference image (derived from a text prompt), our model synthesizes appearance (color image), geometry (scene coordinate map), and semantic (semantic segmentation map) from arbitrary viewpoints, while preserving spatial consistency across modalities. SpatialGen consistently generates superior results to previous methods in our experiments. We are open-sourcing our data and models to empower the community and advance the field of indoor scene understanding and generation.
@@ 1 Introduction
A major consideration in developing 3D scene generation methods is the trade-off between realism and scene diversity. Procedural modeling methods [33, 55, 9] leverage hand-crafted heuristic rules and geometric constraints in the graphics engines, which produce highly realistic and physically plausible indoor environments. However, these scenes lack diversity. Recent 3D generative methods automatically generate scene layouts [28, 46] or other 3D representations like NeRFs [1] and 3D Gaussians [21]. But these methods exhibit limited layout and appearance realism, primarily due to the scarcity of annotated 3D data. In comparison, image-based methods utilize diffusion models to generate panoramas [47, 39] or multi-view images [11, 45] followed by 3D reconstruction. By leveraging powerful 2D priors, these methods show promise in striking a better balance between realism and scene diversity. Image-based methods, however, face additional challenges in multi-view semantic consistency. While recent video generation methods [56, 26, 34] have improved temporal coherence, synthesizing semantically consistent content when exploring beyond input views remain highly challenging.
@@
To overcome these limitations, we collect a new indoor scene dataset on a much larger scale. Our dataset features 4.7M panoramic images with precise 2D and 3D layout annotations, spanning 57,431 rooms and 12,328 scenes. With this dataset, we take a new approach to 3D scene synthesis by building a scalable multi-view diffusion (MVD) model conditioned on 3D layout priors, which achieves high semantic consistency while maintaining the realism and scene diversity in the results.
@@
  • We introduce a new large-scale dataset featuring over 4.7M panoramic images of 57,431 rooms and precise 2D and 3D layout annotations. This dataset fills a critical gap in 3D scene modeling by providing comprehensive multi-view data with structural annotations.
@@ 3 SpatialGen Dataset
As shown in Figure 2, we create physically plausible camera trajectories that navigate smoothly through each scene while avoiding obstacles. These trajectories are sampled at $0.5\textrm{m}$ intervals to ensure comprehensive spatial coverage. For each viewpoint, we generate photorealistic panoramic renderings using an industry-leading rendering engine, capturing color, depth, normal, semantic, and instance segmentation data. We further convert the panoramic image into multiple perspective images using equilib [14]. To ensure both quality and diversity, we apply rigorous filtering criteria during dataset curation, resulting in 12,328 distinct scenes encompassing 57,431 individual rooms with diverse room types. Each scene is annotated with precise 3D layouts and divided into 57,381/50 scenes for training/testing, respectively.
Figure 2 also shows some panoramas and 3D layout annotation from our dataset. Our dataset offers comprehensive structural layout annotations, including architecture elements (i.e., walls, doors, and windows). We further simulate diverse camera motion patterns from panoramic video data to train and evaluate the generation capabilities of existing approaches – a crucial advantage over limited rule-based trajectories from existing dataset. The empirical benefits of this dataset are demonstrated in the experiments.
@@ 5.2.1 Comparison to Score Distillation Methods
As one can see, our method performs slightly worse than SceneCraft when trained solely on Hypersim. We hypothesize that Hypersim is too small for powerful latent multi-view diffusion models like the one employed by our method. When trained on both Hypersim and our datasets, our method outperforms both SDS methods on all metrics. SpatialGen achieves a significantly higher image-reward score than models trained solely on Hypersim, validating the benefit of our large-scale dataset for high-quality 3D scene generation. Furthermore, when tested on our dataset, SpatialGen consistently outperforms the baselines, with a significantly higher image-reward score.
@@
The advantage of our method can be better observed in Figure 6. On both Hypersim and our datasets, SpatialGen generates photorealistic scenes with superior details that are well-aligned with the specified layout. In contrast, SceneCraft struggles to balance layout adherence with text-prompt fidelity, and Set-the-Scene takes nearly two hours to synthesize a radiance field that still lacks fine-grained details. While SpatialGen produces blurry and artifact images, further underscores the benefit of our proposed dataset to the final output quality.
@@ 5.2.2 Comparison to Panorama-as-Proxy Methods
Next, we compare our method to panorama-as-proxy methods. We choose Ctrl-Room [8] as the baseline because the source code and checkpoints are available. We conduct experiments on both Structured3D and our datasets. For Structured3D dataset, we use ground-truth layouts to ensure consistent inputs for both methods. For our dataset, we use the Inward Orbit camera trajectory for consistency.
@@
5.3 More Experiments and Ablations
Additionally, we evaluate SpatialGen on image-to-3D scene generation to validate its generative capability and 3D consistency. Recognizing that well-defined 3D semantic layouts are often unavailable in practice, we further extend the evaluation to video inputs. In this setting, we leverage the state-of-the-art layout estimation model [27] to parse a 3D layout directly from the video. Additional results, implementation details, and ablations are provided in Appendix.B and Appendix.C.
@@ 6 Conclusion & Limitations
We present SpatialGen, a novel framework for layout-guided 3D indoor scene synthesis. At the core of our pipeline is a multi-view multi-modal diffusion model, which generates images with high visual quality and geometric consistency. To train this model, we collect a new synthetic indoor scene dataset with 4.7M panoramic renderings of 57,431 rooms and the 3D layout annotations. These advancements open new possibilities for downstream applications such as interior design, embodied AI, and AR/VR.
@@
First, the cross-view and cross-modal attention introduces additional computational cost to the diffusion model, which limits SpatialGen to generate more images at a time. Second, the camera sampling strategy might affect the generation quality. Moreover, since our method relies on an initial RGB image, it is sensitive to any misalignment between the image and the provided 3D layout. We plan to address these challenges in the future.
@@ Acknowledgments
This work was supported in part by the Key R&D Program of Zhejiang Province (2025C01001) and HKUST Project 24251090T019. We thank Yingqi Shen, Liangbin Hu, and Fuchun Dong (Manycore Tech) for dataset support; Chenfeng Hou for SDS-based experiments; Zhiwei Wang for ControlNet testing; and Kunming Luo for figure suggestions.
Supplementary Material
Figure 8: Examples of SpatialGen dataset.
In the appendix, we provide more details of SpatialGen dataset in Appendix A, additional experimental results in Appendix B, and ablation studies in Appendix C.
@@ A.1 Dataset Construction
The final dataset consists of 12,328 distinct scenes, 57,431 individual rooms covering a variety of room types, and 4.7M photo-realistic panoramic renderings. The total floor area across all scenes is approximately $914,687\textrm{m}^{2}$.
@@ A.3 Dataset Visualization
Appendix B Additional Experiments and Results
In this section, we show more results of text-to-3D scene generation in Section B.1, and conduct comprehensive experiments of image-to-3D scene generation in Section B.2. Furthermore, we show generation results from common videos (unposed) captured in indoor scenes in Section B.3.
@@ Appendix B Additional Experiments and Results
B.1 Text to 3D Scene Generation
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
Forward
SpatialGen (w/o layout) 11.47 0.49 0.59 67.96
SpatialGen (w/ layout) 17.59 0.69 0.32 34.98
Inward orbit
SpatialGen (w/o layout) 12.57 0.48 0.54 64.14
SpatialGen (w/ layout) 17.30 0.66 0.33 35.57
Outward orbit
SpatialGen (w/o layout) 11.14 0.60 0.47 76.73
SpatialGen (w/ layout) 13.32 0.59 0.46 57.76
Random walk
SpatialGen (w/o layout) 11.26 0.45 0.59 98.42
SpatialGen (w/ layout) 14.07 0.62 0.45 52.10
B.2 Image to 3D Scene Generation
Quantitative Results
Qualitative Results
@@ B.2 Image to 3D Scene Generation
In Figure 18, given a reference image (highlighted in orange box) across four diverse camera trajectories , our method successfully generates 3D-consistent novel views and synthesizes semantically plausible content for areas beyond the original input view.
B.3 Video-to-3D Scene Generation
Given the fact that a well-defined 3D layout is not easy to obtain, we apply SpatialGen to the challenging task of generating novel 3D scenes from videos. By leveraging a state-of-the-art layout estimation model, SpatialLM [27], we get the reconstructed 3D layout from the video. Then, we perform text-to-3D scene generation conditioned on this layout and additional user-provided text prompts. This approach allows us to generate entirely new scenes that preserve the structural layout of the original video while altering its stylistic and semantic content based on the text description. We validate this video-to-new-scenes application on the SpatialLM test set. Figure 23 shows qualitative results.
B.4 Failure Cases
Because SpatialGen relies on an initial RGB image, if the images happen to be inconsistent with the given 3D layout, then it fails to generate consistent results. In this example of text-to-3D scene generation in Figure 19, we first convert the provided layout and textual description into an initial RGB image. However, the initialization process yields an RGB output (highlighted in red) that does not align with the given layout—specifically, the orientation of the bed is entirely reversed compared to the ground truth. Consequently, during the subsequent generation stage, the resulting outputs exhibit inconsistencies and fail to align properly with both the layout and the initial RGB image.
Figure 19: Failure case analysis. The initial RGB image, generated from a layout and text, misorients the bed (red, vs. ground truth). This initial error causes subsequent generations to conflict with the layout.
Appendix C Ablation Studies
In this section, we conduct additional ablation studies to verify the layout guidance, design choice of network architecture, and the number of input views.
@@ Appendix C Ablation Studies
Figure 22: Ablation on the effectiveness of the Multi-view Multi-modal Alternating Attention mechanism.
@@
Ablation on Alternating Attention mechanism
We further evaluate the design choice of Multi-view Multi-modal Alternating Attention (AA). We compare our full model (denoted as W/ AA) against a variant that disables the multi-modal attention (denoted as W/O AA). Both models share the same training protocol as the aforementioned.
Qualitative results in Figure 22 reveal a critical weakness in the ablated model. As indicated by the red circles, the W/O AA model fails to produce semantically meaningful segmentations. This deficiency corrupts the geometry in the scene coordinate maps and resulting in less coherent multi-view RGB generation. This demonstrates that without an explicit mechanism to align and refine information across modalities (RGB, geometry, semantics), the model cannot effectively leverage their synergies.
Our complete model W/ AA directly addresses this limitation. By facilitating cross-modal interaction, the AA mechanism enables more precise semantic labels and geometrically consistent scene coordinates. This improvement subsequently elevates the fidelity and view-consistency of the final RGB outputs, confirming that the alternating attention is pivotal for high-quality, multi-modal scene generation.
Ablation on number of input views